# Quantifying Uncertainty: All We Need is the Bootstrap?

### v1

@@ 3.4 Absolute distance from exact CIs
Table 6 shows that B-n has lowest distance from exact, while DB and BCa perform relatively poorly across all groups. This result cannot be interpreted in isolation, because there is typically a trade-off between coverage and distance (see Figure 1 for an illustrative example).
@@
Table 6: Mean absolute distance from exact for one-sided CIs. For each combination, the value is normalized with two standard deviations of exact intervals. The overall column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined.
overall 4 8 16 32 64 128 256 corr mean median Q(0.05) Q(0.95) std
B-n $\underline{0.245}$ $0.423$ $\underline{0.337}$ $\underline{0.276}$ $\underline{0.228}$ $\underline{0.187}$ $\underline{0.151}$ $\underline{0.123}$ $\underline{0.095}$ $\underline{0.140}$ $\underline{0.201}$ $\underline{0.329}$ $\underline{0.330}$ $0.244$
BB $0.271$ $\underline{0.392}$ $0.354$ $0.328$ $0.270$ $0.213$ $0.185$ $0.160$ $\underline{0.095}$ $0.141$ $0.262$ $0.362$ $0.383$ $\underline{0.227}$
BC $0.322$ $0.566$ $0.436$ $0.373$ $0.316$ $0.254$ $0.210$ $0.177$ $0.136$ $0.161$ $0.286$ $0.416$ $0.536$ $0.258$
SB $0.326$ $0.546$ $0.446$ $0.362$ $0.306$ $0.252$ $0.207$ $0.172$ $0.125$ $0.148$ $0.276$ $0.479$ $0.441$ $0.308$
BCa $0.327$ $0.542$ $0.427$ $0.379$ $0.326$ $0.274$ $0.224$ $0.186$ $0.136$ $0.180$ $0.286$ $0.432$ $0.521$ $0.268$
DB $0.353$ $0.557$ $0.488$ $0.387$ $0.331$ $0.290$ $0.240$ $0.193$ $0.159$ $0.224$ $0.299$ $0.488$ $0.501$ $0.279$
PB $0.359$ $0.644$ $0.508$ $0.399$ $0.324$ $0.258$ $0.213$ $0.178$ $0.124$ $0.150$ $0.295$ $0.554$ $0.502$ $0.320$
B-t $1.018$ $4.647$ $0.702$ $0.700$ $0.550$ $0.314$ $0.242$ $0.197$ $0.116$ $0.776$ $1.751$ $0.912$ $1.127$ $0.636$
@@
Table 7 shows where a baseline method has better distance from exact than DB, but only for cases where its coverage is not an order of magnitude worse. We included B-n as the bootstrap method that performs best in distance from exact. Results are similar to those in Table 5 - baseline methods and B-n outperform DB for the two extreme percentiles. That is, in most cases, lower distance from exact comes at the expense of worse coverage.
@@
Table 7: Cases where baseline (with B-n) outperforms DB in absolute distance from exact intervals. The number in the parentheses is the number of combinations where a method outperforms. Note that there are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6.
n functional baseline $\gg$ DB
4 corr fisher (6)
4 mean B-n (9); t-test (9); wilcoxon (9)
4 median B-n (5); m-j (5); q-nonpar (4); q-par (5); wilcoxon (5)
4 Q(0.05) B-n (27); q-nonpar (27); q-par (24)
4 Q(0.95) B-n (21); q-nonpar (10); q-par (21)
4 std B-n (14); chi-sq (14)
8 corr fisher (6)
8 mean B-n (3); t-test (3); wilcoxon (3)
8 median B-n (1); m-j (1); q-nonpar (1); q-par (1); wilcoxon (1)
8 Q(0.05) B-n (30); q-nonpar (28); q-par (24)
8 Q(0.95) B-n (27); q-nonpar (15); q-par (23)
8 std B-n (7); chi-sq (7)
16 Q(0.05) B-n (35); q-nonpar (35); q-par (27)
16 Q(0.95) B-n (16); m-j (16); q-nonpar (2); q-par (16)
16 std B-n (2); chi-sq (2)
32 Q(0.05) B-n (29); m-j (29); q-nonpar (26); q-par (15)
32 Q(0.95) B-n (30); m-j (30); q-nonpar (13); q-par (20)
64 Q(0.05) B-n (8); m-j (8); q-nonpar (8); q-par (4)
64 std B-n (1); chi-sq (1)
128 Q(0.05) B-n (2); m-j (2); q-nonpar (2); q-par (1)
@@
4 Discussion
@@
DB has two weaknesses. It can perform relatively poorly when $n=4,8$ and on extreme percentiles for $n\leq 32$. The latter can be mitigated by using B-n but also raises the question if DB can be modified to deal with these cases, while still preserving most of its simplicity. This leaves us with $n=4,8$, where we have to acknowledge that a non-parametric approach is often worse than a parametric approach even when the assumptions of the parametric method are violated. However, we would argue that these are extremely small and of little relevance to most practitioners. We did not investigate to what extent bootstrap diagnostics (see [6] and [10, Ch.3.10]) could further mitigate these issues.
@@
{acks}
[Acknowledgments] We thank Gregor Sočan for his helpful comments.
@@
Erik Štrumbelj was funded by the Slovenian Research and Innovation Agency research core funding P2-0442.

### v2

@@ 3.4 Absolute distance from exact CIs
Table 3.4 shows that B-n has lowest distance from exact, while DB and BCa perform relatively poorly across all groups. This result cannot be interpreted in isolation, because there is typically a trade-off between coverage and distance (see Figure 1 for an illustrative example).
[…]
@@
Mean absolute distance from exact for one-sided CIs. For each combination, the value is normalized with two standard deviations of exact intervals. The all column is across all experiments, while the remaining results are grouped by sample size or statistical functional. The best performing method for each column is underlined. all 4 8 16 32 64 128 256 corr mean Q0.5 Q0.05 Q0.95 std B-n $\underline{0.245}$ $0.423$ $\underline{0.337}$ $\underline{0.276}$ $\underline{0.228}$ $\underline{0.187}$ $\underline{0.151}$ $\underline{0.123}$ $\underline{0.095}$ $\underline{0.140}$ $\underline{0.201}$ $\underline{0.329}$ $\underline{0.330}$ $0.244$ $B$ $0.271$ $\underline{0.392}$ $0.354$ $0.328$ $0.270$ $0.213$ $0.185$ $0.160$ $\underline{0.095}$ $0.141$ $0.262$ $0.362$ $0.383$ $\underline{0.227}$ BC $0.322$ $0.566$ $0.436$ $0.373$ $0.316$ $0.254$ $0.210$ $0.177$ $0.136$ $0.161$ $0.286$ $0.416$ $0.536$ $0.258$ SB $0.326$ $0.546$ $0.446$ $0.362$ $0.306$ $0.252$ $0.207$ $0.172$ $0.125$ $0.148$ $0.276$ $0.479$ $0.441$ $0.308$ BCa $0.327$ $0.542$ $0.427$ $0.379$ $0.326$ $0.274$ $0.224$ $0.186$ $0.136$ $0.180$ $0.286$ $0.432$ $0.521$ $0.268$ DB $0.353$ $0.557$ $0.488$ $0.387$ $0.331$ $0.290$ $0.240$ $0.193$ $0.159$ $0.224$ $0.299$ $0.488$ $0.501$ $0.279$ PB $0.359$ $0.644$ $0.508$ $0.399$ $0.324$ $0.258$ $0.213$ $0.178$ $0.124$ $0.150$ $0.295$ $0.554$ $0.502$ $0.320$ B-t $1.018$ $4.647$ $0.702$ $0.700$ $0.550$ $0.314$ $0.242$ $0.197$ $0.116$ $0.776$ $1.751$ $0.912$ $1.127$ $0.636$
@@
Table 3.4 shows where a baseline method has better distance from exact than DB, but only for cases where its coverage is not an order of magnitude worse. We included B-n as the bootstrap method that performs best in distance from exact. Results are similar to those in Table 3.3 - baseline methods and B-n outperform DB for the two extreme percentiles. That is, in most cases, lower distance from exact comes at the expense of worse coverage.
[…]
@@
Cases where baseline (with B-n) outperforms DB in absolute distance from exact intervals. The number in the parentheses is the number of combinations where a method outperforms. Note that there are a total of 36 combinations for every pair of $n$ and functional, except Pearson correlation, where there are 6. n functional baseline $\gg$ DB 4 corr fisher (6) 4 mean B-n (9); t-test (9); wilcoxon (9) 4 median B-n (5); m-j (5); q-nonpar (4); q-par (5); wilcoxon (5) 4 Q(0.05) B-n (27); q-nonpar (27); q-par (24) 4 Q(0.95) B-n (21); q-nonpar (10); q-par (21) 4 std B-n (14); chi-sq (14) 8 corr fisher (6) 8 mean B-n (3); t-test (3); wilcoxon (3) 8 median B-n (1); m-j (1); q-nonpar (1); q-par (1); wilcoxon (1) 8 Q(0.05) B-n (30); q-nonpar (28); q-par (24) 8 Q(0.95) B-n (27); q-nonpar (15); q-par (23) 8 std B-n (7); chi-sq (7) 16 Q(0.05) B-n (35); q-nonpar (35); q-par (27) 16 Q(0.95) B-n (16); m-j (16); q-nonpar (2); q-par (16) 16 std B-n (2); chi-sq (2) 32 Q(0.05) B-n (29); m-j (29); q-nonpar (26); q-par (15) 32 Q(0.95) B-n (30); m-j (30); q-nonpar (13); q-par (20) 64 Q(0.05) B-n (8); m-j (8); q-nonpar (8); q-par (4) 64 std B-n (1); chi-sq (1) 128 Q(0.05) B-n (2); m-j (2); q-nonpar (2); q-par (1)
@@
Figure 2: A visualization of the visualization tool for browsing the results of our simulation study. The user can filter results by method and facet based on the dimensions of the experiment (functional, confidence level, distribution).
4 Conclusion
This paper presents the most comprehensive review of empirical results for bootstrap methods to date, along with an extensive empirical comparison. Our study encompasses not only the most widely used non-parametric bootstrap methods but also the most common techniques for quantifying uncertainty in general. Furthermore, we introduce a novel criterion for assessing confidence interval quality, which improves on existing evaluation approaches.
@@ 4 Conclusion
DB has two weaknesses. It can perform relatively poorly when $n=4,8$ and on extreme percentiles for $n\leq 32$. The latter can be mitigated by using B-n but also raises the question if DB can be modified to deal with these cases, while still preserving most of its simplicity. This leaves us with $n=4,8$, where we have to acknowledge that a non-parametric approach is often worse than a parametric approach even when the assumptions of the parametric method are violated. However, we would argue that these are extremely small and of little relevance to most practitioners. We did not investigate to what extent bootstrap diagnostics (see [50] and [3, Ch.3.10]) could further mitigate these issues.
@@
Acknowledgment
We thank Gregor Sočan for his helpful comments.
@@ Acknowledgment
This work was supported by the Slovenian Research and Innovation Agency under Grant P2-0442.
Data Availability Statement
The authors confirm that the data supporting the findings of this study are available within the article and its supplementary materials.